OpenStudy (kimo_indie):

Solve the system of linear equations below. 2x + 3y = 2 x + 6y = 4

OpenStudy (anonymous):

we will mark equation 1 as E1 and equation 2 as E2 first we will switch E1 with E2 2x + 3y = 2 x + 6y = 4 ---> x + 6y = 4 2x + 3y = 2 now will do following E2=E2 - 2*E1 ----> x + 6y =4 ----> x + 6y = 4 (2-2)x + (3-2*6)y =2 -2*4 0*x + -9y = -6 now we will work only in the second equation -9y =-6 y=6/9 =2/3 we will insert it in 2x + 3y = 2 2x +3 *(2/3) =2 2x +2 =2 /-2 2x=0 x=0 the solution will be (0,2/3)